OpenStudy (anonymous):

5 trigonometric functions for sin theta = -4/5;3pi/2

OpenStudy (anonymous):



OpenStudy (anonymous):

find the trig ratios for that triangle

OpenStudy (anonymous):

since a^2 + b^2 = c^2 the side adjacent to the angle must be 3 since cos(x) = adj/hyp tan(x) = opp/adj csc(x) = hyp/opp sec(x) = adj/opp cot(x) = adj/opp ....

OpenStudy (anonymous):

Thanks again, you're a life savor! @Gralwyn

OpenStudy (anonymous):

np